# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VII-B Proof of Theorem 1
$\Phi_{t}\textstyle=\frac{b_{5}(5+\sigma_{g,1}^{2}+L_{g}^{2})\lambda_{x}^{0}}{m(t+1)^{\beta_{1}+v_{x}}}\!+\!\frac{b_{6}^{\prime}\lambda_{x}^{0}}{m(t+1)^{\beta_{3}+v_{x}}}\!+\!\frac{c_{8}(5+\sigma_{g,1}^{2}+L_{g}^{2})+2^{2\varsigma_{x}}d_{1}L_{F}^{2}}{(t+1)^{2\varsigma_{x}}}.$ (159)
Summing both sides of (158) from 00 to $T$ and using the relationship $F(\boldsymbol{x}^{*})\leq F(\boldsymbol{\bar{x}}^{T+1})$, we obtain
$\sum_{t=0}^{T}\!\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}]\!\leq\!2m\mathbb{E}[F(\boldsymbol{\bar{x}}^{0})-F(\boldsymbol{x}^{*})]+2m\sum_{t=0}^{T}\Phi_{t}.$
Combining the preceding inequality and the relationship $\|\nabla F(\boldsymbol{x}^{t})\|^{2}\leq 2\|\nabla F(\boldsymbol{x}^{t})-\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}+2\|\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}$ yields
$\textstyle\textstyle\sum_{t=0}^{T}\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\boldsymbol{x}^{t})\|^{2}]\leq 4m\mathbb{E}[F(\boldsymbol{\bar{x}}^{0})-F(\boldsymbol{x}^{*})]$ (160)
$\textstyle\!+\!4m\sum_{t=0}^{T}\Phi_{t}\!+\!2\sum_{t=0}^{T}\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\boldsymbol{x}^{t})\!-\!\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}].$
By using (141), (159), and (50), the second and third terms on the right hand side of (160) satisfy
$\textstyle 4m\sum_{t=0}^{T}\Phi_{t}\!+\!2\sum_{t=0}^{T}\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\boldsymbol{x}^{t})\!-\!\nabla F(\boldsymbol{\bar{x}}^{t})\|^{2}]$
$\textstyle\leq\frac{2\lambda_{x}^{0}L_{F}C_{x}(v_{x}+2\varsigma_{x})}{v_{x}+2\varsigma_{x}-1}+\frac{4b_{5}(5+\sigma_{g,1}^{2}+L_{g}^{2})\lambda_{x}^{0}(1+v_{x})}{v_{x}}$
@@
$\textstyle\quad+\frac{4mc_{8}(5+\sigma_{g,1}^{2}+L_{g}^{2})2\varsigma_{x}}{2\varsigma_{x}-1}+\frac{m2^{2\varsigma_{x}}d_{1}L_{F}^{2}\varsigma_{x}+2}{2\varsigma_{x}-1}\triangleq c_{9}.$ (161)
We define $c_{10}=4m\mathbb{E}[F(\boldsymbol{\bar{x}}^{0})-F(\boldsymbol{x}^{*})]$ and substitute (161) into (160) to obtain
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[\|\nabla F(\boldsymbol{x}^{t})\|^{2}]\leq\frac{c_{9}+c_{10}}{\lambda_{x}^{0}(T+1)^{1-v_{x}}},$ (162)
which implies (6).
@@ VII-C Proof of Corollary 2
By applying the following inequality:
[…]
to (20), we can obtain
$\epsilon_{i}\leq\frac{\sqrt{2}C_{x}}{\sigma_{i,x}(v_{x}-v_{z}-\varsigma_{i,x})}+\frac{\sqrt{2}C_{y}}{\sigma_{i,x}(v_{y}-\varsigma_{i,y})}+\frac{\sqrt{2}C_{z}}{\sigma_{i,x}(v_{z}-\varsigma_{i,z})}.$ (163)
We denote the given cumulative privacy budget as $\hat{\epsilon}_{i}>0$. According to (163), we can choose DP-noise parameters as $\sigma_{i,x}\!=\!\frac{3\sqrt{2}C_{x}}{(v_{x}-v_{z}-\varsigma_{i,x})\hat{\epsilon}_{i}}$, $\sigma_{i,y}\!=\!\frac{3\sqrt{2}C_{y}}{(v_{y}-\varsigma_{i,y})\hat{\epsilon}_{i}}$, and $\sigma_{i,z}\!=\!\frac{3\sqrt{2}C_{y}}{(v_{z}-\varsigma_{i,z})\hat{\epsilon}_{i}}$, where the parameters $v_{x}$, $v_{y}$, $v_{z}$, $\varsigma_{x}$, $\varsigma_{y}$, and $\varsigma_{z}$ are predetermined constants (e.g., parameters given in Corollary 1) that satisfy the conditions given in Theorem 2. It is clear that a smaller $\hat{\epsilon}_{i}$ leads to larger values of $\sigma_{i,x}$, $\sigma_{i,y}$, and $\sigma_{i,z}$.
Next, we analyze the convergence rate of Algorithm 1 under a strongly convex $F(x)$. According to (134), we have $\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{x}^{*}\|^{2}]\leq\frac{2C_{x}+4mc_{3}c_{4}(\lambda_{x}^{0}\mu^{-1})}{(t+1)^{\beta}}$. The definitions of $C_{x}$, $c_{3}$, and $c_{4}$ (which are given in (50), (132), and the paragraph above (133), respectively) imply that they are positively correlated with the DP-noise parameters $\sigma_{x}^{2}$, $\sigma_{y}^{2}$, and $\sigma_{z}^{2}$, which are all inversely proportional to $\hat{\epsilon}_{i}^{2}$. Therefore, we obtain $\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{x}^{*}\|^{2}]\leq\mathcal{O}(\frac{T^{-\beta}}{\min_{i\in[m]}\{\hat{\epsilon}_{i}^{2}\}})$.
Similarly, according to (146) and (162), the convergence rates of Algorithm 1 under a convex and nonconvex $F(\boldsymbol{x})$ satisfies $\frac{c_{6}+c_{7}}{2\lambda_{x}^{0}(T+1)^{1-v_{x}}}$ and $\frac{c_{9}+c_{10}}{\lambda_{x}^{0}(T+1)^{1-v_{x}}}$, respectively. Given that $c_{6}+c_{7}$ and $c_{9}+c_{10}$ are both positively correlated with DP-noise parameters $\sigma_{x}^{2}$, $\sigma_{y}^{2}$, and $\sigma_{z}^{2}$ which are all inversely proportional to $\hat{\epsilon}_{i}^{2}$, we arrive at the result in Corollary 2.

### v3

@@ VII-B Proof of Theorem 1
$\Phi_{t}\textstyle=\frac{b_{5}(5+\sigma_{g,1}^{2}+L_{g}^{2})\lambda_{x}^{0}}{m(t+1)^{\beta_{1}+v_{x}}}\!+\!\frac{b_{6}^{\prime}\lambda_{x}^{0}}{m(t+1)^{\beta_{3}+v_{x}}}\!+\!\frac{c_{9}(5+\sigma_{g,1}^{2}+L_{g}^{2})+2^{2\varsigma_{x}}d_{1}L_{F}^{2}}{(t+1)^{2\varsigma_{x}}}.$ (152)
Summing both sides of (151) from 0 to $T$ and using the relationship $F(\boldsymbol{x}^{*})\leq F(\mathbf{\bar{x}}^{T+1})$, we obtain
$\sum_{t=0}^{T}\!\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\mathbf{\bar{x}}^{t})\|^{2}]\!\leq\!2m\mathbb{E}[F(\mathbf{\bar{x}}^{0})-F(\boldsymbol{x}^{*})]+2m\sum_{t=0}^{T}\Phi_{t}.$
Combining the preceding inequality and the relationship $\|\nabla F(\mathbf{x}_{i}^{t})\|^{2}\leq 2\|\nabla F(\mathbf{x}_{i}^{t})-\nabla F(\mathbf{\bar{x}}^{t})\|^{2}+2\|\nabla F(\mathbf{\bar{x}}^{t})\|^{2}$ yields
$\textstyle\textstyle\sum_{t=0}^{T}\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\mathbf{x}_{i}^{t})\|^{2}]\leq 4m\mathbb{E}[F(\mathbf{\bar{x}}^{0})-F(\boldsymbol{x}^{*})]$ (153)
$\textstyle\!+\!4m\sum_{t=0}^{T}\Phi_{t}\!+\!2\sum_{t=0}^{T}\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\mathbf{x}_{i}^{t})\!-\!\nabla F(\mathbf{\bar{x}}^{t})\|^{2}].$
By using (44) and (152) the second and third terms on the right hand side of (153) satisfy
$\textstyle 4m\sum_{t=0}^{T}\Phi_{t}\!+\!2\sum_{t=0}^{T}\lambda_{x}^{t}\mathbb{E}[\|\nabla F(\mathbf{x}_{i}^{t})\!-\!\nabla F(\mathbf{\bar{x}}^{t})\|^{2}]$
$\textstyle\leq\frac{2\lambda_{x}^{0}L_{F}\bar{C}_{x}(v_{x}+2\varsigma_{x})}{v_{x}+2\varsigma_{x}-1}+\frac{4b_{5}(5+\sigma_{g,1}^{2}+L_{g}^{2})\lambda_{x}^{0}(1+v_{x})}{v_{x}}$
@@
$\textstyle\quad+\frac{4mc_{9}(5+\sigma_{g,1}^{2}+L_{g}^{2})2\varsigma_{x}}{2\varsigma_{x}-1}+\frac{m2^{2\varsigma_{x}}d_{1}L_{F}^{2}\varsigma_{x}+2}{2\varsigma_{x}-1}\triangleq c_{10}.$ (154)
We define $c_{11}=4m\mathbb{E}[F(\mathbf{\bar{x}}^{0})-F(\boldsymbol{x}^{*})]$ and substitute (154) into (153) to obtain $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[\|\nabla F(\mathbf{x}_{i}^{t})\|^{2}]\leq\frac{c_{10}+c_{11}}{\lambda_{x}^{0}(T+1)^{1-v_{x}}}$.
@@ VII-C Proof of Corollary 2
For any $i\in[m]$, the per-iteration computational cost of Algorithm 1 at iteration $t$ is $(2r+n_{i}r+n_{i})(t+1)$. Hence, the computational complexity of Algorithm 1 over $T$ iterations is $\frac{(2r+n_{i}r+n_{i})(T-1)T}{2}\approx\mathcal{O}(n_{i}rT^{2})$. Furthermore, for any $i\in[m]$, the per-iteration communication cost of Algorithm 1 is $(2r+n)d_{i}$. Hence, the communication complexity of Algorithm 1 over $T$ iterations is $\mathcal{O}((2r+n)d_{i}T)$.
VII-D Proof of Corollary 3
By applying the inequality $\sum_{t=1}^{T}\frac{1}{(t+1)^{r}}\leq\int_{0}^{T}\frac{1}{(t+1)^{r}}dx=\frac{1}{1-r}((T+1)^{1-r}-1)$ to (21), we can obtain
$\epsilon_{i}\leq\frac{\sqrt{2}C_{i,x}}{\sigma_{i,x}(v_{x}-v_{z}-\varsigma_{i,x})}+\frac{\sqrt{2}C_{i,y}}{\sigma_{i,x}(v_{y}-\varsigma_{i,y})}+\frac{\sqrt{2}C_{i,z}}{\sigma_{i,x}(v_{z}-\varsigma_{i,z})},$ (155)
where $C_{i,x}$, $C_{i,y}$, and $C_{i,z}$ are given in (19), (17), and (20), respectively.
We denote the given cumulative privacy budget as $\hat{\epsilon}_{i}>0$. According to (155), we can choose DP-noise parameters as $\sigma_{i,x}\!=\!\frac{3\sqrt{2}C_{i,x}}{(v_{x}-v_{z}-\varsigma_{i,x})\hat{\epsilon}_{i}}$, $\sigma_{i,y}\!=\!\frac{3\sqrt{2}C_{i,y}}{(v_{y}-\varsigma_{i,y})\hat{\epsilon}_{i}}$, and $\sigma_{i,z}\!=\!\frac{3\sqrt{2}C_{i,z}}{(v_{z}-\varsigma_{i,z})\hat{\epsilon}_{i}}$. It is clear that a smaller $\hat{\epsilon}_{i}$ leads to larger $\sigma_{i,x}$, $\sigma_{i,y}$, and $\sigma_{i,z}$.
Next, we analyze the convergence rate of Algorithm 1 under a strongly convex $F(x)$. According to (128), we have $\mathbb{E}[\|\mathbf{x}_{i}^{t}-\boldsymbol{x}^{*}\|^{2}]\leq\frac{2\bar{C}_{x}+4mc_{3}c_{4}(\lambda_{x}^{0}\mu^{-1})}{(t+1)^{\beta}}$. The definitions of $\bar{C}_{x}$, $c_{3}$, and $c_{4}$ (which are given in (44), (126), and the paragraph above (127), respectively) imply that they are positively correlated with the DP-noise parameters $\sigma_{i,x}^{2}$, $\sigma_{i,y}^{2}$, and $\sigma_{i,z}^{2}$, which are all inversely proportional to $\hat{\epsilon}_{i}^{2}$. Therefore, we obtain $\mathbb{E}[\|\mathbf{x}_{i}^{t}-\boldsymbol{x}^{*}\|^{2}]\leq\mathcal{O}(\frac{T^{-\beta}}{\hat{\epsilon}_{i}^{2}})$.
Similarly, the convergence rates of Algorithm 1 under a convex and nonconvex $F(\boldsymbol{x})$ satisfies $\frac{c_{8}}{2\lambda_{x}^{0}(T+1)^{1-v_{x}}}$ and $\frac{c_{10}+c_{11}}{\lambda_{x}^{0}(T+1)^{1-v_{x}}}$, respectively. Given that $c_{8}$ and $c_{10}+c_{11}$ are both positively correlated with DP-noise parameters $\sigma_{x}^{2}$, $\sigma_{y}^{2}$, and $\sigma_{z}^{2}$ which are all inversely proportional to $\hat{\epsilon}_{i}^{2}$, we arrive at the result in Corollary 3.